OpenStudy (anonymous):

Write a recursive rule for this sequence: 4, 6, 9, 13, 18...

OpenStudy (campbell_st):

looks like they increase by 2, 3, 4, 5,....

OpenStudy (zzr0ck3r):

18 is the 5th term 18 = 13+5 hope that helps

OpenStudy (anonymous):

@zzr0ck3r is 18 = 13+5 the recursive rule

OpenStudy (campbell_st):

not quite... but its a big hint

OpenStudy (campbell_st):

a recursive formula uses the previous term to get the new term... its normally has som operation associated with it

OpenStudy (anonymous):

@campbell_st I am so confused

OpenStudy (zzr0ck3r):

ok if we had \[a_1 = 2\\and\\a_2=5 \]then we defined the recursive sequence to be \[a_{n+1} = a_n+3\] and we wanted to know what a_3 was then \[a_3 = a_2+3\\a_3 = 5+3\]

OpenStudy (zzr0ck3r):

now back to your problem notice that third term = the 2nd term + 3 fourth term = the third term + 4 fifth term = fourth term + 5 what does the sixth term =?

OpenStudy (zzr0ck3r):

@sweetpink365 making more sense?

OpenStudy (zzr0ck3r):

well if you come back \[a_{n+1}=a_n+n\] this is the answer. I don't like just giving answers but I want you to understand this. Please message me when you get back on so we can make sure you understand why.

OpenStudy (zzr0ck3r):

well if you come back \[let \space a_n = 4\\then\\a_n\space is\space given\space by\\a_{n+1}=a_n+n\] this is the answer. I don't like just giving answers but I want you to understand this. Please message me when you get back on so we can make sure you understand why